Marina has a total of 21 coins in dimes and quarters. She has two times as many dimes as quarters
a) define the variable
b) state the system of equations that model this situation
c) determine how many dime and quarters marina has
d) how much money does marina have

Answer by stanbon(75887)

Marina has a total of 21 coins in dimes and quarters. She has two times as many dimes as quarters
a) define the variables
d = # of dimes ; q = # of quarters
-------
b) state the system of equations that model this situation
Quantity Eq:: d + q = 21
Quantity Eq:: d = 2q
-----------------------------
c) determine how many dime and quarters marina has
Substitute for "d" and solve for "q"::
2q + q = 21
3q = 21
q = 7 (# of quarters)
----
Solve for "d"::
d = 21-q = 21-7 = 14 (# of dimes)
----------------------------------------
d) how much money does marina have
Ans: 7*25 + 14*10 = 175 + 140 = $3.15
------------
Cheers,
Stan H.
----------------